OpenStudy (anonymous):

Who can help me to solve integration of cos^2 X in term of X?

OpenStudy (anonymous):

\[\int\limits_{}^{}\cos ^{2}x dx\]

OpenStudy (anonymous):

satellit stop taking all the questions im gonna answer

OpenStudy (anonymous):

standard gimmick is to rewrite \[\cos^2(x)=\frac{1}{2}\cos(2x)+\frac{1}{2}\] by one of those long forgotten "power reduction" formulas

OpenStudy (anonymous):

oh sorry i will be quiet

OpenStudy (anonymous):

standard gimmick is to rewrite cos2(x)=12cos(2x)+12 by one of those long forgotten "power reduction" formulas

OpenStudy (anonymous):

top copying me satellite

OpenStudy (anonymous):

i need a medal first